# VDGD: Mitigating LVLM Hallucinations in Cognitive Prompts by Bridging the Visual Perception Gap

### v1

@@ Abstract
Recent interest in Large Vision-Language Models (LVLMs) for practical applications is moderated by the significant challenge of hallucination or the inconsistency between the factual information and the generated text. In this paper, we first perform an in-depth analysis of hallucinations and discover several novel insights about how and when LVLMs hallucinate. From our analysis, we show that: (1) The community’s efforts have been primarily targeted towards reducing hallucinations related to visual recognition (VR) prompts (e.g., prompts that only require describing the image), thereby ignoring hallucinations for cognitive prompts (e.g., prompts that require additional skills like reasoning on contents of the image). (2) LVLMs lack visual perception, i.e., they can see but not necessarily understand or perceive the input image. We analyze responses to cognitive prompts and show that LVLMs hallucinate due to a perception gap: although LVLMs accurately recognize visual elements in the input image and possess sufficient cognitive skills, they struggle to respond accurately and hallucinate. To overcome this shortcoming, we propose Visual Description Grounded Decoding (VDGD), a simple, robust, and training-free method for alleviating hallucinations. Specifically, we first describe the image and add it as a prefix to the instruction. Next, during auto-regressive decoding, we sample from the plausible candidates according to their KL-Divergence (KLD) to the description, where lower KLD is given higher preference. Experimental results on several benchmarks and LVLMs show that VDGD improves significantly over other baselines in reducing hallucinations. We also propose VaLLu, a benchmark for the comprehensive evaluation of the cognitive capabilities of LVLMs.
@@ 1 Introduction
Large Language Models (LLMs) pre-trained at on web-scale text data with the next token prediction objective implicitly compress world knowledge in their parameters zhao2023survey . These models learn general-purpose representations, which can then be aligned with the desired response characteristics zhang2023instruction . This step generally involves fine-tuning on instruction-response pairs, also known as instruction tuning (IT) wei2022finetuned , and is followed by an optional step of reinforcement learning from human feedback (RLHF) bai2022training . Such aligned LLMs can be engaged through text-based inputs, or prompts, directing them to perform various tasks, including those based on information retrieval and reasoning.
Recent advancements in the research community have demonstrated success in multi-modal alignment, whereby fine-tuning LLMs with multi-modal instruction-response pairs enables them to execute tasks that conform to the text prompt as well as additional multi-modal inputs yin2023survey . However, similar to text-only LLMs, Multi-modal Large Language Models (MLLMs) are also prone to hallucination liu2024survey , or the misalignment between factual visual content and corresponding textual responses generated. This is a prominent concern that impacts the reliability and applicability of MLLMs across domains. In this paper, we investigate hallucinations in Large Vision Language Models (LVLMs) to explore the underlying causes and mechanisms.
The research community has made significant attempts to reduce LVLM hallucinations liu2024survey . These attempts include better architectures liu2023improved ; wang2023cogvlm , better alignment techniques chen2024far , and explicit hallucination mitigation techniques yin2023woodpecker ; zhang2024debiasing ; wang2024mitigating . Each approach has demonstrated superior performance over others on benchmark datasets. In this paper, we conduct a comprehensive analysis of LVLM hallucinations to explore the mechanisms and conditions under which LVLMs hallucinate and to measure the community’s progress in addressing these issues. Our results reveal the following:
@@
    Prior work has majorly focused on alleviating hallucinations for visual recognition (VR) prompts (e.g., object hallucinations in prompts that demand image descriptions). While this has led to significant improvements in VR benchmarks (like AMBER), improvements in cognitive prompts that require additional skills, like reasoning and information seeking, have remained stagnant. Additionally, we perform an in-depth analysis to show that there has been relatively less progress in alleviating hallucinations for certain aspects (like descriptions of non-real world scenes) and categories (described in Section 4.2) of VR hallucinations.
@@
    For cognitive prompts we perform controlled experiments to show that LVLMs suffer from a visual perception gap: although they can accurately recognize visual elements in the input image and possess sufficient knowledge and reasoning skills to provide a factual response, they fail to fully perceive the image to generate a factual response.
@@
    Inspired by our findings, we propose a Visual Description Grounding Decoding (VDGD), a simple and novel training-free technique to alleviate LVLM hallucinations. Specifically, we first generate a description of the input image. Next, at every decoding step during auto-regressive response generation, we calculate the Kullback–Leibler divergence (KLD) between the top-k tokens in the current logit and all the logits of the image description. Finally, to predict the next word, we sample from the plausible candidates according to their KLD to the description, where lower KLD is given higher preference. VDGD improves performance on standard benchmarks over other mitigation techniques from literature.
2 Experimental Setup
LVLMs. For the scope of this work, we experiment with 6 different types of LVLMs, namely LLaVA-v1 liu2023visual , LLaVA-v1.5 liu2023improved , LLaVA-v1.6 liu2024llavanext , mPLUG-Owl2 ye2023mplugowl2 , InternLM-X zhang2023internlm and CogVLM wang2023cogvlm , each utilizing a 7B parameter language model, though the number of vision parameters varies. To accurately evaluate their capabilities in visual recognition and perception, we ensure that all models use the same foundational language model. More details can be found in Appendix F.
Evaluation Setup. For evaluating LVLMs, we perform a mix of expert human evaluation and automatic evaluation using GPT-4 Turbo (gpt-4-turbo-2024-04-09). Recent research indicates that using ChatGPT and GPT-4 to score and assess outputs from LLMs aligns well with human evaluations, offering the added benefit of cost reduction zhou2023lima ; ghosh2024closer ; lin2023unlocking ; chan2024chateval ; xu-etal-2023-instructscore ; zheng2023judging . We devise an explainable and multi-aspect evaluation prompt lin2023unlocking ; ghosh2024closer , which prompts the judge (GPT-4) to assign a score between 1 and 5 to an LVLM’s response to an instruction on five aspects: helpfulness, clarity, factuality, depth, and engagement. It also provides an explanation for each score and extracts hallucinated phrases. Prompt is in Appendix D. With hallucinations as our primary focus, all results shown in the main paper are for the factuality metric. The metric is designed to penalize hallucinations and evaluate correctness in open-ended generations. Scores for all metrics and all graphs are in Appendix I.
Evaluation Benchmarks. We employ a range of benchmarks for LVLM evaluation that assess the diverse capabilities of a model. These include AMBER wang2023llm (visual recognition), SynthDoG kim2022donut (OCR), Oven hu2023open (information seeking), MMMU yue2023mmmu (general and holistic reasoning), MMC liu2023mmc , MathVista lu2023mathvista and MATH-Vision wang2024measuring (mathematical reasoning), HallusionBench guan2023hallusionbench , and MME fu2023mme (chart understanding). We only use the sub-sets of these datasets that demand open-ended generation for response, as LLMs have been shown not to be robust to MCQ zheng2024large . Additionally, since automatic evaluation on all these datasets can be expensive, we propose VaLLu, detailed in Section 6.2.
3 Current Progress on Solving LVLM Hallucinations
@@
As LVLMs advance and become more accessible, their application extends beyond merely generating image descriptions. Thus, to better understand hallucinations, we break down information processing by an LVLM, when prompted with an input image and a text instruction, into its constituent steps:
@@
As depicted in Fig. 1, diverse prompts applied to a single image may evaluate the LVLM for various skills. Visual Recognition (VR) involves recognizing visual elements within the image (like objects, backgrounds, etc.), their interrelations, and their unique characteristics.
@@
A prompt that just seeks a description of the image or details about a specific element explicitly evaluates only visual recognition capabilities. Visual Perception (VP), which follows and moves beyond VR, entails interpreting and contextualizing the observed elements in the context of the entire image fu2024blink . For tasks that extend beyond basic image description, accurate visual perception is crucial, surpassing the mere recognition of individual elements. If the prompt seeks further knowledge-specific information related to the elements in the image, then the next step involves extracting relevant information (IE) from the knowledge learned during pre-training. Finally, some prompts may necessitate the LVLM to reason over the visual data, textual prompt, and any implicitly extracted information to deduce a response. While VR relies on the performance of the vision encoder, the abilities to extract accurate information and reason are based on the strength of the foundational language model used in alignment tuning (as alignment does not teach a model new knowledge or skills ghosh2024closer ; zhou2023lima ). VP, which is done implicitly by the model, is a skill taught during alignment tuning and serves as a bridge between VR and other cognitive abilities. barrow1978recovering ; torralba2002depth ; hartley2003multiple
A closer look at the progress in mitigating hallucinations. Fig. 2 (left) compares
@@
the performance of LVLMs across 5 benchmarks. As we clearly see, while newer models that were trained using improved architectures and alignment training data boosted performance on the AMBER benchmark (for recognition), performance on other reasoning and information-seeking benchmarks remained stagnant.
Fig. 2 (right) compares the performance of LLaVA-1.5 across the same 5 benchmarks when employed with explicit hallucination mitigation techniques, namely, Visual Contrastive Decoding (VCD) damonlpsg2023vcd , OPERA huang2023opera , Woodpecker yin2023woodpecker and LURE zhou2023analyzing . As we clearly see, while all these methods boost performance on AMBER, similar to the previous experiments, performance on other reasoning and information-seeking benchmarks remains stagnant.
{mdframed}
[linewidth=1pt, linecolor=black, leftmargin=1pt, rightmargin=1pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=4pt, innerbottommargin=2pt, backgroundcolor=gray!20, roundcorner=5pt] Key Takeaway: Current advancements in reducing hallucinations via improved model architectures, training data, and explicit mitigation strategies are limited to improving the visual recognition capabilities of LVLMs. Progress in addressing hallucinations beyond those related to visual recognition has been relatively quite less.
4 A Closer Look at Visual Recognition Hallucinations
As we have seen in the earlier section, advancements in LVLMs and hallucination mitigation techniques have led to considerable improvements in visual recognition capabilities by these models. In the next few subsections, we take a closer look at visual recognition hallucinations to show that the models and methods discussed above only alleviate hallucinations in some specific cases.
4.1 Moving Beyond Real-World Scenes
As LVLMs are now being employed on a wide range of tasks, considerable research efforts are being made to evaluate the efficacy of LVLMs to reason on images beyond real world
@@ 4.1 Moving Beyond Real-World Scenes
scenes, i.e., charts liu2023mmc or math problems lu2024mathvista . However, how often do LVLMs hallucinate while describing images beyond real-world scenes? Fig 3 (left) compares the performance of various LVLMs across 5 benchmarks. While MMMU and MathVista include images with mathematical figures, MMC includes images with charts from diverse sources, and SynthDoG prompts the model for OCR. For AMBER, MMMU, MathVista, and MMC, we prompt the LVLM with the prompt “Describe the image.” and for SynthDoG we prompt the LVLM with the prompt Extract text from scanned PDFs, photos, and even handwriting.. As we clearly see, compared to AMBER, models hallucinate more often when describing visuals beyond real-world scenes. Fig 3 (right) further shows that hallucination mitigation techniques improve performance on AMBER but rarely other benchmarks.
4.2 Causal Categorization of Visual Recognition Hallucinations
Overview. In our detailed analysis of model responses on the AMBER dataset, we identified intriguing patterns in which hallucinations occur. Although the community has primarily focused on the broader topic of object hallucinations liu2024survey , our findings reveal that visual recognition hallucinations can be classified into several subcategories based on their specific occurrence patterns.
@@ 4.2 Causal Categorization of Visual Recognition Hallucinations
(1) Language Hallucinations: These are hallucinations that are caused by language priors and are unshifted tokens (tokens that would have been output by the non-aligned LLM knowledge through prior context without an image). During auto-regressive decoding, the model might over-rely on the prior tokens in the response rather than the input image. Thus, this may leads the model to mention details in the response from its pre-trained knowledge that is not actually faithful to the image.
(2) Vision Hallucinations: Unlike language hallucinations, the tokens for vision hallucinations are generated by attending to the input image, and are shifted tokens. These are hallucinations caused when the model fails to accurately recognize specific components within an input image, or just inaccurate reasoning or IE, among other factors.
(3) Style Hallucinations. These are hallucinations caused by style imitation. Open-source VLLMs are often trained on instruction-tuning data synthesized from proprietary closed-source models. Training on such datasets makes the VLLM learn to mimic the characteristics of responses in the IT dataset ghosh2024closer ; gudibande2024the (e.g., lengthy answers, a summary line at the end of the response, etc.). However, owing to the lack of sufficient knowledge to mimic its closed-source counterparts, open-source VLLMs might hallucinate facts while responding ghosh2024closer .
@@
(4) Instruction Tuning (IT) Hallucinations. These are hallucinations caused by biases learned during the instruction tuning stage. LLMs are known to mimic the responses associated with the unfamiliar examples in the model’s instruction-tuning data (i.e., those unfamiliar to the pre-trained base model) ghosh2024closer . Formally, this can be defined as distribution shift. As visual instruction tuning is an extreme case of such a distribution shift, IT hallucinations are a major cause of hallucinations in LLMs.
Performance Comparison. Fig. 5 (top;left) compares the frequency of different categories of hallucinations, across various LVLMs for AMBER. Fig. 5 (top;right) displays the frequency of different hallucination categories for LLaVA-v1.5 alongside various hallucination mitigation techniques. Fig. 5 (bottom) offers a similar comparison for MMMU. Our key findings include: (1) Model-based improvements and mitigation strategies have significantly reduced language and vision hallucinations, but progress in style and IT hallucinations remains limited. (2) Specific mitigation techniques are more effective for certain types of hallucinations. For instance, decoding-based methods like VCD and OPERA excel at reducing language hallucinations, whereas object-grounding approaches like Woodpecker are more effective for vision hallucinations. Decoding-based techniques benefit from low-confidence hallucinated tokens and uncertainty encoded in logits as shown in Fig. 4. Further details can be found in Appendix H.1. (3) Existing methods fail to reduce hallucinations on non-VI benchmarks like MMMU. We attribute this to the algorithmic biases of these methods to alleviate VR hallucinations (for example, these algorithms are crafted for real-world scenes).
@@
[linewidth=1pt, linecolor=black, leftmargin=1pt, rightmargin=1pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=4pt, innerbottommargin=2pt, backgroundcolor=gray!20, roundcorner=5pt] Key Takeaway: While community efforts have enhanced the visual recognition capabilities of LVLMs, substantial challenges persist in mitigating hallucinations for images that depict scenarios beyond real-world scenes. Additionally, hallucinations can be distinctly categorized based on their occurrence patterns, and some categories have received lesser attention.
5 The Visual Perception Gap
As illustrated in the previous section, current progress in LVLMs primarily alleviates VR hallucinations. In this section, we delve into LVLM hallucinations in various benchmarks, suggesting that the issue may stem from alignment problems. We first show that for cognitive prompts, LVLMs do not tend to pay attention to the input image but respond only using language priors. Next, we show that this problem may arise from a visual perception gap, i.e., although LVLMs may accurately recognize visual elements in the input image and possess sufficient knowledge and reasoning skills to provide a factual response, they still fail to respond accurately to the prompt and hallucinate.
5.1 Assessing the Extent to Which LVLMs Attend to the Input Image
Figure 6: Base Rank Comparison between AMBER and MATH-Vision datasets as a function of token position in responses generated by CogVLM.
To assess the influence of the image on the LVLM response, we compare the Base Rank (defined in Section 4.2) of tokens across datasets. A lower value would signify that the model solely responds using language priors while a higher value would signify that the model pays attention to the image, i.e., it responds with tokens that are different from what the model would have responded with only language priors. Fig. 6 plots the Base Rank of tokens against the token position in response (average across the dataset). As we see, the Base Rank for AMBER, which prompts for an image description, is higher than MMU, which requires it to reason or extract knowledge based on the image. This hints towards an alignment issue: different from AMBER, for MMMU LVLMs are unable to perceive the image before responding, under-relying on visual cues and over-relying on language priors for generating responses. In Appendix H.1 we explain why decoding-based mitigation techniques discussed in Section 4.2 do not work for cognitive prompts.

### v2

# Visual Description Grounding Reduces Hallucinations and Boosts Reasoning in LVLMs

@@ Abstract
Large Vision-Language Models (LVLMs) often produce responses that misalign with factual information, a phenomenon known as hallucinations. While hallucinations are well-studied, the exact causes behind them remain underexplored. In this paper, we first investigate the root causes of hallucinations in LVLMs. Our findings reveal that existing mitigation techniques primarily reduce hallucinations for visual recognition prompts—those that require simple descriptions of visual elements—but fail for cognitive prompts that demand deliberate reasoning. We identify the core issue as a lack of true visual perception in LVLMs: although they can accurately recognize visual elements, they struggle to fully interpret these elements in the context of the input prompt and effectively link this recognition to their internal knowledge, which is critical for reasoning. To address this gap, we introduce Visual Description Grounded Decoding (VDGD), a simple, robust, and training-free method designed to enhance visual perception and improve reasoning capabilities in LVLMs. VDGD works by first generating a detailed description of the image and appending it as a prefix to the instruction. During response generation, tokens are sampled based on their KL divergence to the description, favoring candidates with lower divergence. Experimental results on multiple visual reasoning benchmarks and LVLMs demonstrate that VDGD consistently outperforms existing baselines 2% - 33%. Finally, we introduce VaLLu, a benchmark designed for comprehensive evaluation of the cognitive capabilities of LVLMs. Code is available here: https://github.com/Sreyan88/VDGD  [^1] Corresponding Author: sreyang@umd.edu. * denotes equal technical contribution.
@@ 1 Introduction
Large Language Models (LLMs) pre-trained on web-scale text data with the next token prediction objective implicitly compress world knowledge in their parameters (Zhao et al., 2023). These models learn general-purpose representations, which can then be aligned with the desired response characteristics (Zhang et al., 2024b). This step generally involves fine-tuning on instruction-response pairs, also known as instruction tuning (IT) (Wei et al., 2022), and is followed by an optional step of reinforcement learning from human feedback (RLHF) (Bai et al., 2022). Such aligned LLMs can be engaged through text-based inputs, or prompts, directing them to perform various tasks, including those based on information retrieval and reasoning.
Recent advancements in the research community have demonstrated success in multi-modal alignment, whereby fine-tuning LLMs with multi-modal instruction-response pairs enables them to execute tasks that conform to the text prompt as well as additional multi-modal inputs (Yin et al., 2023a). Since their introduction, the research community has extensively evaluated the capabilities of Large Vision-Language Models (LVLMs) across areas such as visual recognition (Wang et al., 2023a), perception (Fu et al., 2024), and reasoning (Yue et al., 2024). Researchers have made significant efforts to enhance these capabilities by refining model architectures and training techniques (Zhang et al., 2024a) or by implementing strategies to mitigate hallucinations  [^2] Hallucination refers to the mismatch between factual content and the model’s generated responses. Mitigation aims to reduce these discrepancies for more accurate outputs and improve task performance. (Huang et al., 2023a).
Among these approaches, training-free hallucination mitigation techniques stand out as a cost-effective and accessible method to enhance LVLM performance without the need for extensive computational resources. However, current hallucination mitigation techniques primarily focus on reducing hallucinations and improving the performance of visual recognition tasks (e.g., describing a scene, OCR, etc). Our extensive experiments show that these techniques fall short when applied to cognitive prompts requiring reasoning or knowledge extraction, indicating a significant gap in their ability to address hallucinations in more complex, reasoning-intensive scenarios.
To this end, we propose Visual Description Grounded Decoding (VDGD), a simple and novel training-free technique designed to improve LVLM reasoning by reducing hallucinations for cognitive prompts. Much like humans, who often write down their observations of an image in their own words and refer back to them while tackling complex reasoning tasks, we hypothesize that grounding the LVLM’s response generation in an explicit visual description can significantly aid its reasoning capabilities. Specifically, we first generate a description of the input image. Next, at every decoding step during auto-regressive response generation, we calculate the Kullback–Leibler divergence (KLD) between the top-k tokens in the current logit and all the logits of the image description. Finally, to predict the next word, we sample from the plausible candidates according to their KLD to the description, where lower KLD is given higher preference. By continuously referencing its descriptive summary of the visual content, the LVLM maintains alignment between the visual input and its generated responses, thereby reducing hallucinations. To summarize, our contributions are as follows:
@@
    We conduct extensive controlled experiments to investigate the causes of hallucinations in LVLMs, categorizing them into distinct types. We demonstrate that existing hallucination mitigation techniques are effective only for visual recognition (VR) prompts—specifically, those that involve object recognition and image descriptions—while proving ineffective for cognitive prompts that require reasoning. We further show that these techniques are biased towards reducing language hallucinations in real-world scenes but fail to address hallucinations in non-real-world contexts and other categories.
@@
    Identifying the core issue, we reveal that LVLMs suffer from a critical visual perception gap. Although these models can recognize visual elements and possess the reasoning skills needed for factual response generation, they struggle to fully perceive the visual information in the context of the prompt, resulting in hallucinated outputs for cognitive reasoning tasks.
@@
    We introduce VDGD, a novel and training-free method that enhances LVLM reasoning by grounding response generation in visual descriptions. We evaluate VDGD on X benchmarks with prompts that require deliberate reasoning and how VDGD significantly outperforms existing techniques, improving performance by 2% - 33% by reducing hallucinations.
  4. 4.
    Finally, we introduce VaLLu, a meticulously curated benchmark designed to evaluate the cognitive capabilities of LVLMs. Comprising 1,500 carefully selected instances, VaLLu is constructed by refining existing benchmarks and rigorously removing noisy and misleading examples identified through our thorough analysis. VaLLu emphasizes a balanced mix of diverse cognitive tasks, focusing exclusively on questions that require open-ended responses.
2 Related Work
Large Vision-Language Models In recent years, Large Vision-Language Models (LVLMs) have seen rapid advancements with the release of numerous new models. These developments have been driven by innovations in model architectures (Liu et al., 2023b; Ye et al., 2023; Wang et al., 2023b), training methods (Liu et al., 2023b), alignment techniques (Chen et al., 2024b), and data augmentation techniques (Liu et al., 2023b; Wang et al., 2024b). To assess the capabilities of these models, the research community has introduced a variety of benchmarks specifically designed to evaluate different aspects, including visual recognition and understanding in diverse scenarios (Liu et al., 2023a; Kim et al., 2022), reasoning (Lu et al., 2023; Yue et al., 2024), knowledge-based information retrieval (Hu et al., 2023), and other specialized tasks.
Reducing Hallucinations to Improve Response Quality: Hallucination mitigation broadly aims to enhance the accuracy of model responses. While numerous studies have proposed methods to reduce hallucinations and improve response accuracy (Sicong Leng et al., 2023; Huang et al., 2023b; Yin et al., 2023b; Zhou et al., 2023b) and evaluated their effectiveness (Wang et al., 2023a; Guan et al., 2023), most focus on mitigating object hallucinations in visual recognition tasks. While these efforts are valuable, they fall short of assessing the true reasoning capabilities of LVLMs. Moreover, despite the growing focus on evaluation and mitigation, the underlying causes of hallucinations remain largely unexplored.
3 How Close Are We to Reliable LVLM Responses?
Experimental Setup for Analysis. We evaluate six LVLMs—LLaVA-v1, LLaVA-v1.5, LLaVA-v1.6, mPLUG-Owl2, InternLM-X, and CogVLM—all built on a 7B parameter language model. The models are tested across seven benchmarks: AMBER (visual recognition), SynthDoG (OCR), MMMU (expert-level reasoning), MathVista and MATH-Vision (mathematical reasoning), MMC (chart understanding), and MME and HallusionBench. To address potential limitations in existing evaluation metrics, which may not fully capture reasoning skills and often rely on string-matching techniques, we employ a combination of expert human evaluation and LLM-as-a-Judge (using GPT-4-turbo-2024-04-09). This paradigm has been extensively followed by prior studies in LLM evaluation (Zhou et al., 2023a; Ghosh et al., 2024; Zheng et al., 2023) and hallucination evaluation (Sicong Leng et al., 2023; Yu et al., 2023; Liu et al., 2023c). We devise an evaluation prompt that penalizes hallucinations and assigns scores from 1 to 5 based on factual correctness (see Appendix G). This approach also helps us standardize scoring across benchmarks and has a higher correlation ($\approx$0.97%) to expert human evaluations (see Section D). Quantitative results for all graphs in the paper are in Appendix L.
@@ 3 How Close Are We to Reliable LVLM Responses?
Preliminaries. As LVLMs advance and become more accessible, their application extends beyond merely generating image descriptions. Thus, to better understand visual perception and reasoning, we break down information processing by an LVLM into its constituent steps:
@@
[…]
[trim= 0cm 2.25cm 0cm 1cm, width=]perception.pdf
@@
Figure 1 demonstrates that different prompts applied to a single image can assess various LVLM skills. Visual Recognition (VR) focuses on identifying visual elements and their relationships within the image, such as describing objects or specific details. Visual Perception (VP) extends beyond VR by interpreting and contextualizing these elements within the broader scene (Fu et al., 2024), essential for tasks requiring more than basic recognition. Prompts that demand knowledge-specific insights (also known as information-seeking prompts) engage in knowledge extraction (KE) learned from pre-training. Finally, prompts requiring reasoning involve combining visual data, textual prompts, and extracted knowledge to generate a response. While recognition depends on the vision encoder, knowledge extraction, and reasoning rely on the foundational language model. VP, implicitly learned during alignment (Ghosh et al., 2024; Zhou et al., 2023a), bridges VR and higher cognitive functions, facilitating comprehensive understanding of the image in context to the prompt (Barrow et al., 1978; Torralba & Oliva, 2002; Hartley & Zisserman, 2003).
3.1 Do Hallucination Mitigation Techniques improve Cognitive Abilities?
[…]
[trim= 0cm 0.75cm 0cm 0.5cm, width=]radar1-2.pdf
@@ 3.1 Do Hallucination Mitigation Techniques improve Cognitive Abilities?
Fig. 2 (left) compares the performance of LVLMs across 5 benchmarks. As we clearly see, while newer models that were trained using improved architectures and alignment training data boosted performance on the AMBER benchmark (for recognition), performance on other reasoning and information-seeking benchmarks remained stagnant.
Fig. 2 (right) compares the performance of LLaVA-1.5 across the same 5 benchmarks when employed with explicit hallucination mitigation techniques, with strategies ranging from improving model architectures to explicit response correction strategies, namely, Visual Contrastive Decoding (VCD) (Sicong Leng et al., 2023), OPERA (Huang et al., 2023b), Woodpecker (Yin et al., 2023b) and LURE (Zhou et al., 2023b). As we clearly see, while all these methods boost performance on AMBER, performance on other reasoning and information-seeking benchmarks remains stagnant.
3.2 Do Mitigation Techniques Generalize Beyond Real-World Scenes?
[…]
[trim= 0cm 0.75cm 0cm 0.5cm, width=]radar3-4.pdf
@@ 3.2 Do Mitigation Techniques Generalize Beyond Real-World Scenes?
As LVLMs are now being employed on a wide range of tasks, considerable research efforts are being made to evaluate the efficacy of LVLMs to reason on images beyond real-world scenes, i.e., charts (Liu et al., 2023a) or math problems (Lu et al., 2024). However, how often do LVLMs hallucinate while describing images beyond real-world scenes? Fig 3 (left) compares the performance of various LVLMs across 5 benchmarks. MMMU and MathVista include images with mathematical figures and MMC includes charts from diverse sources. On the other hand, SynthDoG prompts the model for OCR and AMBER for image descriptions for real-world scenes. As we clearly see, compared to AMBER, models hallucinate more often when describing visuals beyond real-world scenes. Fig 3 (right) further shows that hallucination mitigation techniques improve performance on AMBER but rarely other benchmarks.
[…]
[width=]vllm.drawio5.pdf
3.3 Are all hallucinations equal?
Overview. In our detailed analysis of model responses on the AMBER dataset, we identified intriguing patterns in which hallucinations occur. Although the community has primarily focused on the broader topic of object hallucinations (Liu et al., 2024b), our findings reveal that visual recognition hallucinations can be classified into several subcategories based on their specific occurrence patterns. Specifically, we perform the token-distribution analysis proposed by  (Lin et al., 2023).
@@ 3.3 Are all hallucinations equal?
(1) Language Hallucinations: These are hallucinations that are caused when LVLMs over-rely on the language priors (or prior tokens in the response) rather than the input image. We attribute all hallucinated tokens that are unshifted as language hallucinations.
[…]
@@
(2) Vision Hallucinations: Unlike language hallucinations, the tokens for vision hallucinations are generated by attending to the input image and are shifted or marginal tokens. These are hallucinations caused when the model fails to accurately recognize visual elements in an input image or inaccurate reasoning or knowledge extraction, among other factors.
(3) Style Hallucinations. These are hallucinations caused by style imitation. Open-source VLLMs are often trained on instruction-tuning data synthesized from proprietary closed-source models. Training on such datasets makes the VLLM learn to mimic the characteristics of responses in the IT dataset (Ghosh et al., 2024; Gudibande et al., 2024) (e.g., lengthy answers, a summary line at the end of the response, etc.). However, owing to the lack of sufficient knowledge to mimic its closed-source counterparts, open-source VLLMs might hallucinate facts while responding (Ghosh et al., 2024).
(4) Instruction Tuning (IT) Hallucinations. These are hallucinations caused by biases learned during the instruction tuning stage. LLMs are known to mimic the responses associated with the unfamiliar examples in the model’s instruction-tuning data (i.e., those unfamiliar to the pre-trained base model) (Ghosh et al., 2024). Formally, this can be defined as a distribution shift. As visual instruction tuning is an extreme case of such a distribution shift, IT hallucinations are a major cause of hallucinations in LLMs.
Performance Comparison. Fig. 5 (top;left) compares the frequency of different categories of hallucinations across various LVLMs for AMBER. Fig. 5 (top; right) displays the frequency of different hallucination categories for LLaVA-v1.5 alongside various hallucination mitigation techniques. Fig. 5 (bottom) offers a similar comparison for MMMU. Our key findings include: (1) Model-based improvements and mitigation strategies reduce language and vision hallucinations, but progress in style and IT hallucinations remains limited. (2) Specific mitigation techniques are more effective for certain types of hallucinations. For e.g., decoding-based methods like VCD and OPERA excel at reducing language hallucinations, whereas grounding approaches like Woodpecker are more effective for vision hallucinations. Decoding-based techniques benefit from low-confidence hallucinated tokens and uncertainty encoded in logits discussed next. More discussion is in Appendix K.1. (3) Existing methods fail to reduce hallucinations on reasoning benchmarks like MMMU. We attribute this to the algorithmic biases of these methods that are crafted to just reduce object hallucinations.
@@
[linewidth=1pt, linecolor=black, leftmargin=1pt, rightmargin=1pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=4pt, innerbottommargin=2pt, backgroundcolor=gray!20, roundcorner=5pt] Key Takeaway: Current hallucination mitigation techniques mainly improve LVLMs’ visual recognition skills but are less effective in improving other cognitive abilities, such as reasoning. Additionally, these improvements are largely limited to real-world scenes and specific types of hallucinations, leaving areas like non-real-world scenes largely unexplored.
4 The Visual Perception Gap: LVLMs can see but not perceive
It is evident that current hallucination mitigation techniques primarily enhance visual recognition but fail to improve other cognitive abilities, such as reasoning or knowledge extraction. In this section, we investigate the underlying cause of this limitation. Our analysis reveals that LVLMs often rely on language priors rather than attending to the input image when generating responses to reasoning prompts. This leads us to identify a critical issue: the visual perception gap. While LVLMs can accurately recognize visual elements and possess the necessary knowledge and reasoning skills to respond factually, they struggle to perceive and interpret these elements in relation to the input prompt. This gap in perception causes hallucinations, resulting in incorrect responses.
4.1 Visual Blind Spots: LVLMs Ignore the Input Image for Reasoning
[…]
[trim= 0.5cm 1.5cm 0cm 1cm, width=]TR_line_cogvlm_first_token.pdf
Figure 6: Base Rank Comparison between AMBER and MATH-Vision datasets as a function of token position in responses (for CogVLM).
With findings from previous sections, we now assess the influence of the image on the LVLM response. Precisely, we compare the Base Rank (defined in Section 3.3) of tokens across datasets. A lower value would signify that the model solely responds using language priors, while a higher value would signify that the model pays attention to the image, i.e., it responds with tokens different from what the model would have responded with only language priors. Fig. 6 plots the Base Rank of tokens against the token position in response (average across the dataset). As we see, the Base Rank for AMBER, which prompts for an image description, is higher than Math-Vision, which requires it to reason or extract knowledge based on the image. This hints towards an alignment issue: different from AMBER, for Math-Vision LVLMs are unable to perceive the image before responding, under-relying on visual cues and over-relying on language priors for generating responses. In Appendix K.1 we explain why decoding-based mitigation techniques discussed in Section 3.3 do not work for reaosning in cognitive prompts.